### v2

# Xinyuan Lu1    Liangming Pan 2    Yubo Ma3    Preslav Nakov4    Min-Yen Kan1

Abstract
Current Large Language Models (LLMs) exhibit limited ability to understand table structures and to apply precise numerical reasoning, which is crucial for tasks such as table question answering (TQA) and table-based fact verification (TFV). To address these challenges, we introduce our Tool-Augmented Reasoning framework for Tables (TART), which integrates LLMs with specialized tools. TART contains three key components: a table formatter to ensure accurate data representation, a tool maker to develop specific computational tools, and an explanation generator to maintain explainability. We also present the ToolTab dataset, a new benchmark designed specifically for training LLMs in table–tool integration. Our experiments indicate that TART achieves substantial improvements over existing methods (e.g., Chain-of-Thought) by improving both the precision of data processing and the clarity of the reasoning process. Notably, TART paired with CodeLlama achieves 90.0% of the accuracy of the closed-sourced LLM GPT-3.5-turbo, highlighting its robustness in diverse real-world scenarios. All the code and data are available at https://github.com/XinyuanLu00/TART.
1 Introduction
Tabular data is prevalent across multiple fields such as scientific research, financial reporting, and healthcare records [11]. Manual handling of such data can be both routine and error-prone, or may require specialized skills, highlighting the need for automated table reasoning [2]. Typical table-based reasoning tasks include table question answering (TQA), which extracts precise information from tables to answer given queries [7, 48, 25], and table-based fact verification (TFV), which verifies the correctness of statements by cross-referencing them with facts stored in tables [33, 22].
Figure 1: Example of the TableQA task, demonstrating the verification of travel time via boat schedule and demonstrating the critical skills needed for accurate table reasoning: table structure understanding, precise numerical calculations, and executing sequential reasoning steps.
Figure 2: An overall framework of TART, which contains three main modules: (itable formatter, which prepares and organizes the raw table data, (iitool maker, which creates specialized tools for precise table manipulation, and (iiiexplanation generator, which produces user-friendly explanations integrating the output of different tools.
Modern large language models (LLMs) such as GPT-4 [26] have shown remarkable reasoning capabilities across a variety of tasks, spurring interest in their application to table-based tasks [41]. However, table-based reasoning presents unique challenges for LLMs, which are primarily trained on sequential text data [44], as illustrated by a real-world example in Figure 1. (1) Understanding and operating on the structure of teh table: LLMs must adapt to the non-linear format of the tables, which demands unique reasoning skills such as recognizing headers, interpreting the roles of the rows and the columns, and precisely extracting information from relevant table cells. (2) Precise numerical reasoning: Tables often contain quantitative information that requires precise calculations and comparisons. LLMs must perform operations such as summation, averaging, or trend analysis accurately, often over multiple cells or tables, which is a shift from their usual text-based reasoning tasks [15, 18]. (3) Reasoning planning: LLMs often need to plan several reasoning steps ahead. This includes decomposing the original question, determining which table parts are relevant and anticipating intermediate calculations or data transformations.
Existing approaches that use LLMs for table-based reasoning can be broadly classified into two categories. One is chain-of-thought (CoT) reasoning [37], in which the model is prompted to perform step-by-step reasoning over the input table flattened as a textual sequence [44, 16, 4, 41]. Despite its flexibility, CoT often lacks precision in tabular operations and numerical reasoning, such as sorting, counting, and filtering [39]. On the other hand, program-based reasoning (PoT) [12, 5] prompts the model to generate SQL or Python code to enable precise reasoning [19, 45, 36, 39]. However, this method tends to struggle with planning and its reasoning is less understandable for humans [45]. Therefore, there is potential value in integrating the advantages of program-based and textual reasoning, to achieve both high precision and explainability in table-based reasoning.
Inspired by the recent paradigm of tool-augmented language models [35, 31], we propose Tool-Augmented Reasoning framework for Tables (TART), which integrates external tool calling into the chain-of-thought reasoning process, as shown in Figure 2. Initially, TART processes the input table using a specialized module table formatter to clean and to format the raw table data, preparing it for the subsequent table operations. Subsequently, the tool maker calls specialized tools (Python functions) for tabular manipulation and numerical reasoning (e.g., adding columns, selecting rows, and grouping). Alongside these tools, TART also crafts a reasoning plan that outlines the programmatic calling sequence of the tools, specifying the necessary arguments and the expected return values for each call. Finally, following the structured reasoning plan, the explanation generator produces a hybrid text-and-program output, integrating calls to external tools into coherent, human-readable chain-of-thought explanations. Hence, TART efficiently delegates table operations and precise numerical calculations to generated tools, while preserving the planning and explainability of CoT.
To train the modules in TART, we further synthesize the ToolTab dataset by distilling knowledge from a teacher LLM. We evaluate TART on nine different table-based reasoning benchmarks. The results highlight the effectiveness of integrating task-specific tools for enhancing complex reasoning capabilities. Notably, TART consistently outperformed the CoT baseline, achieving near-parity with GPT-3.5-turbo on several benchmarks, showcasing its practical utility in real-world scenarios.
2 Related Work
Table-based reasoning tasks involve interpreting and manipulating data from structured tabular sources to answer questions, verify facts, or generate summaries. Early approaches used executable SQL or SPARQL to interact with tabular data [42, 43], or graph neural networks to better encode table structures [46, 40]. However, they typically suffer from poor generalization capabilities due to their reliance on specific table formats and linguistic patterns.
Recently, large language models (LLMs) have shown great potential in improving table-based reasoning tasks. Pioneering studies have focused on developing table pre-training strategies, where LLMs are trained with sentence-table pairs to enhance their general table reasoning capabilities [6, 15, 47, 13, 41]. Subsequent work further explored different reasoning strategies to improve efficacy. For example, Ye et al.[41] proposed using LLMs as decomposers to effectively reason over tables by breaking down large tables into smaller sub-evidence and complex questions into simpler sub-questions. Want et al.[36] introduced the Chain-of-Table framework, where tabular data is explicitly used in the reasoning chain for intermediate thoughts. ReAcTable [45] combined reactive and proactive reasoning strategies to improve accuracy and retrieval from complex tables. Despite these advances, these approaches primarily rely on pure textual reasoning, such as chain-of-thought, which sacrifices the precision required for table manipulations and numerical reasoning, both crucial for table-based reasoning.
To address these limitations, TART extends the use of LLMs with integrated external tools, aimed at improving reasoning precision with function execution while maintaining explainability with LLM-based reasoning. Unlike previous methods that focus on a single reasoning strategy, TART enhances table reasoning with multiple strategies: formatting tables for better data representation, calling tools for precise calculation, and generating user-friendly explanations that clarify the reasoning process.
3 Methodology
Generally, a table-based reasoning model, $f_{\theta}(\cdot)$, parameterized by $\theta$, takes an input query $Q$ and a table $\mathcal{T}$ to produce a response $Y=f_{\theta}(Q,\mathcal{T})$. Based on this generic formulation, the nature of $Q$ and $Y$ differs depending on the specific table reasoning task: in table-based QA, $Q$ is a question and $Y$ is the answer; in table-based fact verification, $Q$ is a claim and $Y$ is its veracity label; in table summarization, $Q$ specifies the requirement and $Y$ is the summary. A table $\mathcal{T}$ is characterized by a caption $P$ and its contents ${T_{i,j}\mid i\leq R_{T},j\leq C_{T}}$, where $R_{T}$ and $C_{T}$ represent the number of rows and columns, respectively. Each cell $(i,j)$ contains data $T_{i,j}$.
[…]
1. Table Formatter. TART first transforms the original table $\mathcal{T}$ with guidance from the query $Q$ into a formatted table $\mathcal{T}^{\prime}$. The formatter optimizes data formats, aligns columns, and adjusts data types as needed for the query, producing a well-formatted table that is used in subsequent reasoning.
2. Tool Maker. Given $\mathcal{T}^{\prime}$ , the tool maker generates a set of candidate tools $S$ useful for solving $Q$. It also develops a reasoning plan $R$ that details the high-level reasoning, which includes the tool calling order, as well as the necessary arguments and the expected return values for the tool calls.
3. Explanation Generator. Given the reasoning plan $R$ as a programmatic guide for chain-of-thought reasoning, the explanation generator is responsible for producing a user-friendly explanation $E$ that incorporates the use of the tools. The explanation also concludes with the final answer $\mathcal{A}$, derived from the execution of the reasoning plan $R$.
3.1 Table Formatter
We first train a specialized open-sourced large language model as the table formatter $\mathcal{F}$, which transforms the noisy raw input table $\mathcal{T}$, into a more structured and manageable format, $\mathcal{T}^{\prime}$, to facilitate subsequent reasoning
$\mathcal{T}^{\prime}=\mathcal{F}(\mathcal{T},Q)$ (1)
where the output table $\mathcal{T}^{\prime}$ is formatted in three aspects. (1) Data Cleaning: the model formats the cell values, such as removing currency symbols and textual footnotes to facilitate the execution of external functions to perform table operations or numerical reasoning. (2) Data Standardization: It converts different data representations into a uniform format, e.g., transforms the data from formats like “MM/DD/YYYY” to a consistent “YYYY-MM-DD” format across the entire table. (3) Error Handling: the model is also responsible for fixing obvious errors or missing values in the table, such as automatically inferring header names for columns without the table headers. We introduce the table formatter to ensure that the data in the input table is uniform and optimized for subsequent reasoning, especially to make it more compatible with function execution. In practice, we transform the formatted table $\mathcal{T}^{\prime}$ into a Python array, facilitating easier interpretation and processing by subsequent reasoning modules.
3.2 Tool Maker
Recent studies have shown that LLMs have the capability of synthesizing relevant tools by understanding the problem context and creating solutions based on the crafted tools [31, 3, 35]. Motivated by this, we train another specialized LLM $\mathcal{M}$ as a tool maker, which takes as input the reformatted table $\mathcal{T}^{\prime}$ and the query $Q$ to generate a set of candidate tools $S$ and develops a reasoning plan $R$ that details the high-level reasoning steps.
$S,R=\mathcal{M}(\mathcal{T^{\prime}},Q).$ (2)
The tool set $S=\{s_{1},\cdots,s_{n}\}$ consists of $n$ specialized tools and each tool $s_{i}$ is a Python function that performs table operations (e.g., get_column_by_name), numerical reasoning (e.g., average, argmax), or higher-level functions (e.g., linear_regression). In table-based reasoning, these automated tools are essential to handle reasoning tasks that textual-based LLMs cannot address effectively, thereby significantly enhancing their problem-solving capacities.
Unlike previous work that manually defined a small number ($<10$) of hand-crafted tools [20, 27] or retrieved tools from a predefined set [29, 23], we choose to train a specialized tool maker model that learns to generate tools dynamically, based on the specifics of the table and the context of the problem. This approach not only preserves the model’s ability to “extract” previously encountered tools from its parametric memory, but also empowers the model to create novel tools as needed for unique problems, as shown in Section 4.4. While generating tools offers greater flexibility, it is crucial to prevent the tool maker from creating overly-specific tools (e.g., count_people_on_third_floor), as this would hinder its ability to generalize to new problems. To address this issue, we incorporate tool abstraction and tool deduplication steps when constructing synthetic data for training the module (elaborated on in Section 3.4).
In addition to generating useful tools, the model also constructs a high-level reasoning plan $R=[r_{1},\cdots,r_{N}]$, which outlines how the tools should be applied. The reasoning plan is formulated as a sequence of $N$ function calls. Each function call $r_{i}=(s_{i},A_{i},V_{i})$ includes the function $s_{i}\in\mathcal{S}$, the argument $A_{i}$ passed to the function, and the variable $V_{i}$ that stores the result of the function call $s_{i}(A_{i})$. This reasoning plan acts as a programmatic blueprint, guiding the table-based reasoning process. Both the tool set $S$ and the reasoning plan $R$ are then provided to the explanation generator, which produces the final explainable reasoning output.
3.3 Explanation Generator
While program-based reasoning plans are precise, they are often difficult for non-expert users to understand. Moreover, certain types of reasoning, such as commonsense or narrative-based reasoning, are better communicated in natural language. To address this, TART incorporates a specialized module called the explanation generator $\mathcal{E}$, which generates chain-of-thought natural language explanations integrated with function calls, following the steps outlined in the reasoning plan $R$
$O=\mathcal{E}(S,R).$ (3)
The final output $O$ of TART provides detailed explanations for the function calls. For example, the function call get_column_by_name is explained as, “First, retrieve the column listing snowfall in inches.” Additionally, the explanation generator groups related function calls together to create coherent and easy-to-follow explanations, as illustrated in Figure 2.
3.4 Model Training
As no prior work adopts the tool-augmented LLM framework for table reasoning, there does not exist training data to train the modules Table Formatter, Tool Maker and Explanation Generator ($\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$) in TART. Previous studies have demonstrated that smaller LLMs can learn from distilling the generated outputs of larger teacher LLMs that have better reasoning capabilities [38, 34, 17]. Following this, we use a teacher LLM $\mathcal{L}$ to first synthesize tool-integrated solution trajectories for a set of seed table-based reasoning tasks. These high-quality solution trajectories serve as the blueprint from which we automatically extract and rearrange their components to build training sets for $\mathcal{F}$, $\mathcal{M}$, and $\mathcal{E}$, the three modules used in TART.
Training Data Synthesis.
For all modules, we use GPT-4 as the teacher LLM $\mathcal{L}$ to generate training data. As shown in Table 4, we select five diverse table reasoning datasets: two from TQA and three from TFV, spanning general knowledge (Wikipedia) as well as domains such as finance, health, and scientific documents. These datasets provide a broad range of reasoning types. We few-shot prompt $\mathcal{L}$ to generate tool-integrated solutions for training instances for each dataset. In each solution, the model is prompted to clean the table, invent useful tools, and propose a reasoning plan along with explanations. We provide our prompt in Appendix F.
After generating the solutions, we evaluate the final answers against the ground truth, retaining only the instances with correct answers. Subsequently, we refine the solutions by removing overly specific tools through tool abstraction and tool deduplication. Tool abstraction filters out tools that appear only once, keeping those with broader applicability. Tool deduplication consolidates similar tools that perform the same function, but have different names or implementations (e.g., add and sum). As a result, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, respectively. We refer to this training dataset as ToolTab, with detailed statistics provided in Table 5.
Training Configurations.
Instruction fine-tuning [24, 10] has emerged as a critical strategy that directs LLMs to adhere to specified instructions, facilitating their reasoning capability across a wide range of table-based tasks. Therefore, we use open-source LLMs with instruction tuning as the backbone models for the modules of TART, specifically Llama2-7B [32], Llama3-8B, CodeLlama-7B [30] and Deepseek-Coder-7B-Instruct-V1.5 [14]. We fine-tune all TART modules independently on their respective training datasets from ToolTab, using the standard next-token prediction objective.
TableFV TableQA
Model Setting TabFact PubHT SCITAB TabMWP FinQA Avg. Acc.
I. TableLlama w/o Fine-tuning 72.3 72.5 67.4 46.8 3.2 52.4
w/ DirectQA 72.9 70.5 74.2 48.4 3.7 54.0
II. Llama2-7b w/ DirectQA 64.4 81.2 64.0 55.3 6.4 54.3
w/ CoT 52.6 55.0 42.7 74.5 4.2 45.8
w/ TART 69.2 55.0 53.4 88.8 19.2 57.1 (+24.7%)
[…]
w/ CoT 48.4 62.4 41.0 88.3 8.5 49.7
w/ TART 69.7 68.5 47.2 92.6 27.1 61.0 (+22.7%)
[…]
w/ CoT 45.2 51.7 38.8 70.7 2.7 41.8
w/ TART 66.5 69.8 44.9 90.1 25.0 59.3 (+41.9%)
[…]
w/ CoT 52.1 62.4 45.5 84.6 8.5 50.6
w/ TART 71.3 69.1 47.8 93.1 30.9 62.4 (+23.3%)
III. GPT-3.5-turbo w/ TART 78.7 63.6 59.3 88.3 56.4 69.3
GPT-4 w/ TART 87.7 84.1 63.6 98.3 68.5 80.4
Table 1: Performance evaluation across backbone models using the TART framework, highlighting the best (bold) and second-best (underlined) results. The accuracy is calculated on testing sets, with overall average accuracy in the last column (Avg. Acc.). The red number indicates the average increase percentage over the CoT methods.
4 Experiments
Datasets and Baselines
To rigorously evaluate the performance of our proposed TART framework, we select two categories of benchmarks for table-based reasoning. (1) Table question answering (TQA) benchmarks: WikiTableQuestion (WTQ) [28] focuses on simple factoid questions. HiTab (HIT) [9], TabMWP (TMP) [21] and FinQA (FQA) [8] datasets focus on numerical reasoning reasoning. TAT-QA (TAT) [48] and HybridQA (HYQ) [7] require joint reasoning over the table and the text for financial reports and Wikipedia tables, respectively. 2) Table-based fact verification (TFV): We select TabFact (TAF) [6], SCITAB (SCT) [22], and PubHealthTab (PHT) [1] datasets, which focus on verifying facts based on tables from Wikipedia, scientific articles, and public health articles, respectively.
For baseline comparisons, we select well-known table-based open-source LLMs such as TableLlama [44] , as well as text-pretrained models (Llama2-7b, Llama3-8b) and code-pretrained models (CodeLlama-7b, DeepSeek-Coder-7b). We choose the 7b and the 8b versions to represent a balance between computational efficiency and the capacity for complex reasoning and generalization. For each model, we fine-tune with two settings: (1) DirectQA, where models generate answers directly from questions and tables, and (2) Chain-of-Thought (CoT) reasoning, which requires models to formulate a step-by-step reasoning process before concluding with an answer.
Implementation
For TART, we use the answer given by executing the reasoning plan; if the reasoning plan is not executable, we use the answer given by CoT. For each model, we train the model with ToolTab while leaving the rest (WTQ, HIT, TAT, and HYQ) as held-out unseen datasets. All experiments were conducted on a GPU server with Intel Xeon Platinum 8480C (224) @ 2.900GHz CPU and 8 NVIDIA H100 (80G) GPUs. The training process for Llama-2-7b-hf, CodeLlama-7b-hf, and deepseek-coder-7b-instruct-v1.5 requires a single GPU for approximately 20 hours, using a batch size of 4, learning rate of 5e-5, sequence length of 1500, gradient accumulation steps of 2, and 10 training epochs. Training Llama-3-8b required up to two GPUs for around 20 hours with the same settings. To minimize randomness, a temperature of 0.0 was used, while all other hyperparameters for sampling the output from the LLMs remained at their default values. For the closed-source version of TART, we use GPT-3.5-turbo and GPT-4 with two in-context examples.
4.1 Main Results
We first evaluate TART and the baselines on in-domain datasets, where their training sets are used to construct ToolTab. The experimental results, as shown in Table 1, reveal a notable performance improvement in our model compared to baseline models. We have four major observations.
1. TART consistently outperforms CoT across all four backbone models and datasets. For example, with CodeLlama-7b as the backbone model, TART outperforms DirectQA and CoT by 16.3% and 41.9% on average, respectively. This highlights the effectiveness of integrating task-specific tools in enhancing complex reasoning capabilities.
2. With CodeLlama-7b as the backbone model of TART, it achieves the highest accuracy increase of 41.9%., whereas Llama3-8b shows the least improvement of 22.7%. This discrepancy is likely because of CodeLlama-7b’s specialized pre-training in coding tasks, which enhances the capabilities of creating tools for structured queries and operations.
3. The performance gains of TART also vary for different datasets, with FinQA showing the highest increase, while PubHealthTab shows the least. This discrepancy suggests that the financial focus of the FinQA dataset, which demands extensive numerical reasoning and structured data manipulation, benefits significantly from the TART approach.
4. Using closed-source models (GPT-3.5-turbo and GPT-4) as the backbone models for TART achieves an average accuracy of 74.9, significantly outperforming the open-source counterparts, which average at 60.0 accuracy. Nonetheless, the highest-performing open-source model DeepSeek-7b reaches up to 90.0% of GPT-3.5-turbo’s performance and 77.6% of GPT-4, illustrating the competitiveness of open-source models in the creation and use of tools despite the model size gap.
4.2 Out-of-Domain Results
TQA Hybrid TQA
Model Setting HIT WTQ TAT HYQ
Llama2-7b w/ DirectQA 19.1 23.4 15.4 8.5
w/ CoT 22.1 12.7 20.0 6.7
w/ TART 19.2 17.0 17.0 6.4
[…]
w/ CoT 33.8 26.8 29.3 11.0
w/ TART 34.6 32.5 29.3 12.2
[…]
w/ CoT 16.1 22.6 14.6 9.7
w/ TART 22.3 30.3 13.8 9.0
[…]
w/ CoT 20.5 26.2 15.3 8.1
w/ TART 29.8 33.5 17.0 11.2
Table 2: Out-of-Domain evaluation results for TART framework, highlighting the best (bold) and the second-best (underlined) results.
We hypothesize that the TART has enhanced generalization capabilities compared to CoT due to its ability to create and use general tools. To validate this, we further evaluate TART across four different out-of-domain (OOD) datasets: HiTab (HIT), WikiTableQuestion (WTQ), TAT-QA (TAT), and HybridQA (HYQ). The results are shown in Table 2.
The TART method demonstrates variable effectiveness, with notable improvements in certain contexts. For instance, it achieves an average accuracy increase of 29.3% on the WTQ dataset, indicating robust domain-transfer capabilities. The Deepseek-7b backbone model particularly excels, with 30.6% increase in accuracy. We hypothesize that this superiority stems from its pre-training on coding tasks, which equips it with the capability of effectively creating and using tools in novel domains, surpassing pure-text-based pretraining models such as Llama-2-7b. The analysis in Section 4.4 supports our hypothesis, suggesting that TART excels in developing generic table reasoning functions that generalize well across different domains.
Tab Formt TAF PHT SCT TMP FQA
Llama-2 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7
Llama-3 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
Codellama 67.6/78.0 81.2/66.1 64.6/53.9 94.1/91.5 76.1/35.7
DeepSeek 70.7/79.7 72.5/71.3 63.5/51.3 95.7/93.9 74.5/38.6
Tool Mkr TAF PHT SCT TMP FQA
Llama-2 70.2/81.8 65.8/60.2 53.9/61.5 95.7/91.1 61.7/31.0
Llama-3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5
Codellama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8
Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0
Table 3: Results of TART with different backbone modules. The top section uses deepseek-code-7b as the Tool Maker, while the bottom section uses Llama-3-8b as the Table Formatter. The best performance is highlighted in bold and the second-best is underlined. Tab Formt stands for Table Formatter and Tool Mkr for Tool Maker.
4.3 Impact of Foundation Models
To explore the optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 3. We find that using Llama-3-8B as the table formatter and DeepSeek-7B as the tool maker achieves the best average execution rate (76.8) and accuracy (68.6). This aligns with our expectations given that Llama-3-8B excels in processing long tables while DeepSeek-7B, with its pre-training on code, demonstrates superior capability in tool creation. Detailed results are shown in Table 6 in Appendix B.
Rank Tool Name
1 get_column_by_name
2 get_column_cell_value
3 get_row_index_by_value
4 extract_price
5 get_row_by_name
6 equal_to
7 get_column_by_index
8 subtract
9 divide
10 add
Figure 3: Analysis of tool usage in our TART framework. Shown are the distribution (left) and the categories of the top-20 tools across models (center), as well as the top-10 tools in TART Codellama-7b (right).
4.4 Analysis of Tool Creation
We then performed an in-depth analysis of how TART creates and utilizes tools shown in the Section 4.4.1 and Section 4.4.2.
4.4.1 Tool Distribution.
Figure 3 (left) illustrates the tool usage distribution across different backbone models in TART, highlighting a long-tail distribution. The most frequently used tools are primarily associated with table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), aligning with our observations in Section 4.1. Figure 3 (right) provides a detailed breakdown of tool categories for the top-30 tools, showing that table preprocessing and numerical reasoning tools are the most prevalent. This supports the consistency of tool utilization patterns within TART.
4.4.2 Tool Overlap on OOD Datasets.
Figure 5 illustrates the tool overlap between in-domain datasets and OOD datasets. We find that code pre-training models (CodeLlama-7b and DeepSeek-7b) exhibit a tendency to reuse existing tools when adapting to OOD data. However, text pre-training models demonstrate less overlap, indicating that they tend to solve problems by crafting new tools. The tendency to reuse tools might explain why code pre-training models gain better generalization capabilities in unfamiliar data.
Figure 4: Tool overlap distribution between InD and OOD datasets across different models.
Figure 5: Comparison of the number of repeat tools and correct tools across different models.
4.4.3 Tool Creation and Usage Analysis.
Figure 5 further reveals that although Llama2-7b frequently reuses tools, it often applies them inappropriately. In contrast, CodeLlama-7b not only exhibits a high rate of tool reuse, but also demonstrates a greater accuracy in their appropriate application. Meanwhile, Llama3-8b, despite its lower rate of tool reuse, excels in the correct usage of tools, which contributes to its superior performance.
Figure 6: Case study of TART comparing it to CoT reasoning. Panel (a) illustrates a numerical calculation error in CoT where incorrect arithmetic leads to a wrong answer, and panel (b) demonstrates a table location error where CoT fails to retrieve the correct table values. Both errors can be reduced by TART through tool integration.
4.5 Case Study
To gain deeper insights into the advantages of TART over CoT reasoning, we conducted a case study, shown in Figure 6. The examples highlight the limitation of CoT in numerical reasoning and table preprocessing, such as incorrect calculation (Figure 6 a) and incorrect retrieval (Figure 6 b). Conversely, TART overcomes these challenges effectively via integrating specialized tools like subtract and get_column_by_index. Despite these strengths, TART still encounters issues related to data type mismatches and incorrect programming plans. A detailed analysis of error types in TART can be found in Appendix E.
5 Conclusion and Future Work
In this paper, we introduce an open-source framework to improve table-based reasoning through the Tool-Augmented Reasoning framework for Tables (TART). This framework solves the challenges of current LLMs’ limited ability to understand table structure and execute precise numerical calculations, and maintains explainability. TART consists of a table formatter for accurate data representation, a tool maker for creating specialized tools, and an explanation generator maintaining interpretable explanations. To train TART, we present the ToolTab dataset, a novel benchmark containing a diverse set of real-world tables and their tool-augmented solutions. Experiments across nine benchmarks show that integrating our TART method into different open-sourced LLMs enhances accuracy on table-based reasoning. Furthermore, in-depth analysis revealed that TART effectively learns and uses tools. Future work could extend TART to a multimodal framework by incorporating image-based question-answering and fact-verification to generate richer explanations. Additionally, generating explanations to satisfy the needs of different end users, such as laypeople and experts, could further improve the TART’s applicability and impact.
Limitation
Despite the promising results, our proposed framework has certain limitations that warrant further investigation:
Computational Complexity.
The TART model may affect efficacy, especially when handling simple questions in quick-response scenarios.
Dataset Coverage.
While our efforts have focused on expanding the range of our dataset to include a variety of tableQA datasets, some table-related datasets remain unrepresented in ToolTab. As a result, despite TART’s capacity to adapt to different OOD datasets and tasks, its performance might still different with the complexities and unique challenges of new table tasks and datasets that it has not yet encountered. Having initiated the development of an expansive, versatile tool-enhanced model for tables, we encourage continued research in this area to further advance the model’s ability to generalize across diverse table configurations.
Ethic Statement
Transparency and Integrity.
We ensure that all methodologies, data sources, and technologies used in this study are disclosed transparently. We aim to provide a comprehensive and honest account of our findings, acknowledging both the capabilities and limitations of our proposed solution.
Data Privacy and Security.
Our research uses datasets that are either publicly available or were collected with explicit consent. We adhere to strict data privacy and security protocols to protect the information and ensure it is used solely for the purposes of this research.
Acknowledgements
This research is supported by the Ministry of Education, Singapore, under its MOE AcRF TIER 3 Grant (MOE-MOET32022-0001). The computational work for this article was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).
Appendix A Dataset Composition for TART Training
In Table 4, we show the composition of the seed datasets used for training our TART model. These datasets vary in terms of the tasks, the domains, and the types of input and output data. For instance, TabMWP and FinQA focus on TableQA tasks within mathematics and finance domains respectively, requiring a combination of tables, text, and questions as inputs, with short answers as outputs. Meanwhile, PubHealthTab, TabFact, and SCITAB target table fact-checking tasks across health, general Wikipedia, and scientific article domains. These datasets similarly involve tables and claims as inputs but differ in the specifics of the domain-related claims, each producing a short label as an output. To construct the ToolTab, we obtain 11,701, 9,916, and 9,916 training instances for the table formatter $\mathcal{F}$, tool maker $\mathcal{M}$, and explanation generator $\mathcal{E}$, respectively. Detailed statistics are provided in Table 5.
Dataset Task Domain Input Output
1. TabMWP TableQA Maths Table, Question Answer (Short)
2. FinQA TableQA Finance Table, Text, Question Answer (Short)
3. PubHealthTab Table Fact Checking Health Table, Claim Label (Short)
4. TabFact Table Fact Checking Wikipedia Table, Claim Label (Short)
5. SCITAB Table Fact Checking Scientific Articles Table, Claim Label (Short)
Table 4: Statistics about the seed datasets for TART training, highlighting their respective tasks, domains, and the nature of input and output data.
Dataset Train Dev Generated Executable Table Tool Explanation
Sample Sample Formatter Maker Generator
TabMWP 23,059 7,686 6,000 5,835 6,000 5,713 5,713
FinQA 6,251 883 1,984 1,609 1,967 1,148 1,148
TabFact 92,283 12,792 1,866 1,773 1,866 1,701 1,705
PubHealthTab 1,180 152 1,180 1,075 1,180 958 958
SciTab 690 - 690 625 688 396 396
Total 123,463 21,513 5,720 10,917 11,701 9,916 9,916
Table 5: Statistics about dataset ToolTab for training the TART model.
Appendix B Different Backbone Combinations
In the pursuit of identifying optimal module combinations within the TART framework, we explore various pairings of table formatter and toolmaker modules shown in Table 6. The combination of Llama-3-8B as the table formatter and DeepSeek-7B as the tool maker performs the most effective pairing, having the best average execution rate and accuracy (76.8 and 68.6 respectively). This best combination aligns with our expectations given that Llama-3-8B excels in processing long tables while DeepSeek-7B, with its pre-training on code, demonstrates superior capability in tool creation.
Module Name TableFV TableQA Avg.
Table Formatter Tool Maker TabFact PubHealthTab SCITAB TabMWP FinQA Exe./Acc.
Llama-2 Llama-2 64.9/79.5 65.8/59.2 55.1/60.2 90.4/91.8 65.4/26.0 68.3/63.3
Llama-2 Llama-3 70.7/75.9 73.2/65.1 64.0/46.5 91.0/93.6 60.6/37.7 71.9/63.8
Llama-2 Codellama 70.2/76.5 73.8/74.5 64.6/56.5 94.7/88.8 71.8/34.1 75.0/66.1
Llama-2 Deepseek 71.8/78.5 75.8/66.4 64.0/57.0 93.6/92.0 73.4/37.7 75.7/66.3
[…]
Llama-3 Llama-3 75.5/75.4 71.1/69.8 63.5/52.2 97.9/92.4 62.2/38.5 74.0/65.7
Llama-3 Codellama 75.5/85.2 74.5/71.2 62.9/57.1 95.7/91.7 68.1/39.8 75.3/69.0
Llama-3 Deepseek 76.6/84.7 79.2/67.8 62.4/55.9 94.1/94.4 71.8/40.0 76.8/68.6
[…]
CodeLlama Llama-3 66.5/71.2 75.2/69.6 62.4/57.7 94.1/91.0 60.1/36.3 71.2/65.2
CodeLlama Codellama 64.9/75.4 77.9/75.0 68.5/50.8 95.2/92.2 71.3/34.3 75.6/65.5
CodeLlama DeepSeek 67.6/78.0 81.2/66.1 64.6/53.9 94.1/91.5 76.1/35.7 76.7/65.0
[…]
DeepSeek Llama-3 66.5/80.8 65.1/69.1 63.5/54.0 94.1/93.2 59.6/42.9 69.8/68.0
DeepSeek CodeLlama 67.0/80.2 71.1/70.8 58.4/52.9 96.8/90.1 69.7/36.6 72.6/66.1
DeepSeek DeepSeek 70.7/79.7 72.5/71.3 63.5/51.3 95.7/93.9 74.5/38.6 75.4/67.0
Table 6: The TART framework with different backbone modules. The best performance is bold. The second best performance is underlined.
Appendix C Tool Use on Different Backbone Models
Table 7 shows the top-10 tools that dominate the table processing (e.g., get_column_by_name) and numerical reasoning (e.g., add), consistent with our earlier findings in Section 4.1. Further illustrating this, Figure 3 (b) presents a tool categorization for the top 30 functions. We can see that table preprocessing tools constitute the highest percentage at 71.0%, followed by numerical reasoning tools at 21.8%. Together, these categories account for over 90% of tool usage, verifying our assumption that TART is better at table preprocessing and numerical reasoning.
Rank Llama2 Llama3 DeepSeek
1 get_column_by_name get_column_by_name get_column_by_name
2 get_column_cell_value get_column_cell_value get_column_cell_value
3 get_row_index_by_value get_row_index_by_value get_row_index_by_value
4 extract_price extract_price extract_price
5 equal_to equal_to get_row_by_name
6 get_column_by_index get_row_by_name equal_to
7 subtract get_column_by_index divide
8 get_row_by_name divide get_column_by_index
9 add subtract subtract
10 multiply add add
Table 7: Comparison of the top 10 functions across TART-Llama2-7b, TART-Llama3-8b, and TART-DeepSeek-7b
Appendix D CoT Baseline Implementation
For a direct and fair comparison with TART, the same number of CoT samples are generated using the same IDs from the TART training dataset. These samples are generated using GPT-4, prompted with two ICL examples (detailed in Appendix F). In total, we have 9,916 training instances.
Similarly to TART, the CoT baseline was implemented across four different backbone models: Llama2-7b-hf, Llama3-8b, CodeLlama-7b-hf, and DeepSeek-Coder-7b-Instruct-V1.5. Each model was instructed to generate a step-by-step reasoning explanation followed by the final answer as per the instructions: INSTRUCTION: Given the following table, and question, generate a step-by-step reasoning explanation and the final answer.
The training process was aligned with that of TART to ensure experimental consistency. Llama-2-7b-hf, CodeLlama-7b-hf, and deepseek-coder-7b-instruct-v1.5 each requires a single GPU for approximately 12 hours, using a batch size of 4, a learning rate of 5e-5, a sequence length of 1500, gradient accumulation steps of 2, and 10 training epochs. Training Llama-3-8b requires up to 2 GPUs for around 10 hours with the same settings.
Appendix E Error Analysis
To precisely categorize error types in CoT reasoning, we annotate 50 randomly selected CoT error cases. The result (Figure 7) shows that the major error type is incorrect numerical reasoning, followed by errors related to table operations. This analysis verifies the necessity for our proposed TART, which addresses these issues by integrating specialized numerical and table operation tools.
Figure 7: The error types and their distribution of CoT reasoning and TART framework.
Appendix F Prompts
We provide detailed prompts of the TART framework, including the tool discovery process and explanation generation process.
Tool Discovery Prompt:
Task Description: Given a table and a question,
the task is to generate a python program to
answer the question.
Requirements:
1. First define some functions to be used in the program.
2. Try to reuse the functions defined in the previous problems if possible.
3. When defining a new function, make sure this function is general enough to be used in other problems.
4. Define a function called solution(table_data) that takes the table data as input and returns the answer to the question.
Table: Table Content
Question: Question
Answer: Answer
table_data = table data array
#FUNCTION1 Description
def FUNCTION1():
Function Body
#FUNCTION2 Description
def FUNCTION2():
Function Body
def solution(table_data):
Solution Body
return answer
print(solution(table_data))
[[FUNCTION_SOLUTION]]
Explanation Generation Prompt:
Task: Transform Python code used for a table question answering task into an easily understandable explanation in natural language embedded with function calls.
Follow these requirements:
1. The explanation should be the natural language combined with bracketed segments $<$ $>$ for code.
2. The code segments in the brackets $<$ $>$ should indicate the line number of the code, with the format: ###<line number>.
3. Multiple lines of codes are separated with ’;;;’ in the brackets $<$ $>$.
Table: Table Content
Question: Question
Answer: Answer
Python Code:
table_data = table data array
def solution(table_data):
Line 1 ###1
Line 2 ###2
Line 5 ###5
return answer
print(solution(table_data))
Output Explanation:
First, we should get the column
that ... $<$###1 ;;; ###2$>$.
Finally, we find that $<$###5$>$.
[[OUTPUT_EXPLANATION]]
CoT Prompt:
Task Description: Given a table and a question, the task is to generate a step-by-step reasoning explanation and the final answer.
Table: Table Content
Question: Question
Answer: Answer
Python Code:
table_data = table data array
Output Explanation:
To answer this question, first, we ...
Second, to determine..., we compare ...
Therefore, the answer is ...
[[OUTPUT_EXPLANATION]]
TART (GPT-4) Prompt for Table Formatter:
Task Description: Given the following table, context and question, format the table into a python array.
Table: Table Content
Question: Question
Answer: Answer
Python Code:
table_data = table data array
[[LINEARIZED_TABLE]]
TART (GPT-4) Prompt for Tool Maker:
Task Description: Given the following table, context and question, the table_data, generate the python code to solve it.
Table: Table Content
Question: Question
Answer: Answer
table_data = table data array
Python Code:
#FUNCTION1 Description
def FUNCTION1():
Function Body
#FUNCTION2 Description
def FUNCTION2():
Function Body
def solution(table_data):
Solution Body
return answer
print(solution(table_data))
[[FUNCTION_SOLUTION]]